Acetylene, C2H2 (also known as ethyne), is commonl
Answer #1
Concepts and reason

The chemical bond is formed by the attraction of atom or ions. Due to electronegativity, the atoms can attract the electrons in the chemical bond. The chemical bonding is formed with the help of overlap of atomic orbitals.

There are two types of bonds. They are sigma bond (σ)\left( \sigma \right) and pi bond (π)\left( {\rm{\pi }} \right) .

Fundamentals

Sigma bond:

The head – on overlapping of atomic orbitals which leads to a covalent bond formation is represented as sigma bond.

Pi bond:

A side wise overlapping of atomic orbitals which leads to the formation of bond is represented as pi bond.

Example: — сH2
Н.С:

The double bond in ethene contains 1 sigma bond and 1 pi bond. The 4 valance electrons of each carbon forms a bond with 2 hydrogen atom, 1 carbon – carbon sigma bond and 1carbon – carbon pi bond.

sidewise overlap n
н
head on overlapping
Н
head on overlapping

The double bond in ethene includes 1 σ(sp2sp2overlap){\rm{\sigma }}\,{\rm{(s}}{{\rm{p}}^{\rm{2}}}{\rm{ - s}}{{\rm{p}}^{\rm{2}}}{\rm{ overlap)}} and 1 π(ppoverlap){\rm{\pi }}\,{\rm{(p}}\,{\rm{ - }}\,{\rm{p overlap)}} bonds and each carbon – hydrogen bond shows σ(sp2soverlap){\rm{\sigma }}\,{\rm{(s}}{{\rm{p}}^{\rm{2}}}{\rm{ - s overlap)}} .The bond angle is 1200 and shows Trigonal planar geometry.

H CEC
-H

п
п
о
Н
Н

&B
С -Сп

C-C o

С - Но

Ans:

8
С-Со
С - Сп
C-Ησ